# Observer Design for Singularly Perturbed Linear Networked Control Systems Subject to Measurement Noise

### v1

@@ Abstract
thanks: This work was supported by the Australian Research Council under the Discovery Project DP200101303, the France Australia collaboration project IRP-ARS CNRS and the ANR COMMITS ANR-23-CE25-0005, and by ANID through the FONDECYT Initiation grant 11250715. E-mail corresponding author: weixuanw@student.unimelb.edu.au
@@ 1 Introduction
In addition to the challenges introduced by multiple time scales, many practical systems also involve internal states that cannot be directly measured. For example, in battery management, the state-of-charge and state-of-health are not directly observable and must instead be inferred. In such scenarios, observers are indispensable tools for estimating the unmeasurable states. Moreover, electrochemical, thermal and aging dynamics of battery evolve on multiple time scales, which naturally lead to the use of multi-time-scale observers (zou2016multi). Extensive research has been conducted on the multi-time-scale observer design for both continuous-time and sampled-data singularly perturbed systems, e.g., lin2010composite; yang2016observer; ferrante2021observer, while several studies also investigated observers for NCS, e.g., postoyan2011framework; wang2017observer. Despite these advances, multi–time–scale observers for SPNCS that specifically handle measurement noise, which frequently occurs in practice, while accounting for time–scale separation and network phenomena such as aperiodic transmissions and scheduling protocols remain unexplored.
@@ 4 Stability analysis
$|\xi(t,j)|_{\mathcal{E}}\leq k|\xi(0,0)|_{\mathcal{E}}e^{-a(t+j)}+\gamma_{v_{1}}\|v_{1}\|_{(t,j)}\\ +\gamma_{v_{2}}\|v_{2}\|_{(t,j)}+\gamma_{\dot{u}_{s}}\|\dot{u}_{s}\|_{(t,j)},$
@@
The proof is provided in the appendix. It is important to note that our result, presented in Theorem 1, shows that the ultimate bound depends on the derivative of the input signal and the magnitude of the measurement noise. This result is a special case of the standard DISS introduced in DISS.
@@ 5.1 Numerical analysis
Consider plant (1) with $A_{11}=\left[\begin{smallmatrix}a_{1}&0\\ 0&-a_{5}\end{smallmatrix}\right]$, $A_{12}=\left[\begin{smallmatrix}a_{2}&0\\ 0&0\end{smallmatrix}\right]$, $A_{21}=\left[\begin{smallmatrix}0&0\\ a_{3}&0\end{smallmatrix}\right]$, $A_{22}=\left[\begin{smallmatrix}-a_{2}&0\\ -a_{2}&-a_{4}\end{smallmatrix}\right]$, $C_{1s}=\left[\begin{smallmatrix}c_{1}&c_{2}\end{smallmatrix}\right]$, $C_{2s}=\left[\begin{smallmatrix}-1&0\end{smallmatrix}\right]$, $C_{2f}=\left[\begin{smallmatrix}0&-1\end{smallmatrix}\right]$, where $a_{1}=10^{-3}$, $a_{2}=0.34$, $a_{3}=1.1$, $a_{4}=4.9$, $a_{5}=1.2$, $c_{1}=-0.21$ and $c_{2}=1.9$. To simplify the computation, we assume observer gains have the form $L_{1s}=\left[\begin{smallmatrix}n_{1}\\ 0\end{smallmatrix}\right]$, $L_{1f}=\left[\begin{smallmatrix}0\\ a_{6}\end{smallmatrix}\right]$, $L_{2s}=\left[\begin{smallmatrix}-n_{2}\\ -n_{2}\end{smallmatrix}\right]$, $L_{2f}=\left[\begin{smallmatrix}0\\ 0\end{smallmatrix}\right]$.
@@
Since $u_{s}$, $y_{p_{s}}$ and $y_{p_{f}}$ are scalars, the protocols are given by $h_{s}(\kappa_{s},e_{s})=0$ and $h_{f}(\kappa_{f},e_{f})=0$, which are UGES protocols. Define $W_{\star}(\kappa_{\star},e_{\star})\coloneqq|e_{\star}|$, $s\in\{s,f\}$. Then, Assumption 1 holds with $\underline{a}_{W_{\star}}(s)=\overline{a}_{W_{\star}}(s)=s$, $\lambda_{\star}=0$ and $M_{\star}=1$. Moreover, inequality (18) is satisfied with $L_{s}=1.82$ and $A_{H_{s}}=A_{21}^{s}$, and (20) holds with $L_{f}=0$ and $A_{H_{f}}=A_{21}^{f}$.
@@
By solving the LMI (21) and (23) simultaneously while maximizing $\epsilon^{*}T^{*}$, we obtain the parameters in observer gains to be $a_{6}=0.16$, $n_{1}=0.02$ and $n_{2}=0.01$. Meanwhile, we have $P_{s}=\left[\begin{smallmatrix}0.11&-0.62\\ -0.62&3.96\end{smallmatrix}\right]$, $\gamma_{s}=2.5$, $a_{\rho_{s}}=0.13$, $P_{f}=\left[\begin{smallmatrix}3.95&0.14\\ 0.14&2.73\end{smallmatrix}\right]$ and $\gamma_{f}=1.64$, $a_{\rho_{f}}=2.69$. We select $\tau_{\text{mati}}=0.169\ s<T(L_{s},\gamma_{s},\lambda_{s})$. Then, with $\lambda_{s}^{*}=0.33$, $\eta_{s1}=0.1$, $\eta_{s2}=0.1$, $\eta_{s3}=0.5$ and $\eta_{s}=\eta_{s1}+\eta_{s2}+\eta_{s3}=0.7$, Lemma 1 is satisfied (see Section A for details). Furthermore, we pick $\lambda_{f}^{*}=0.456$, then $T^{*}=T(L_{f},\gamma_{f},\lambda_{f}^{*})=0.42\ \nicefrac{{s}}{{\epsilon}}$.
@@
For the given $P_{f}$ and $W_{s}$, the minimum possible values for $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$ and $\lambda_{5}$ in (30) are given by $0.005$, $0.17$, $0.02$, $0.35$ and $0.02$, respectively. Similarly, for the specified $W_{s}$, $P_{s}$, $W_{f}$ and $P_{f}$, it can be shown that $b_{1}=9.1$, $b_{2}=0.19$, $b_{3}=2.8$$a_{\Delta_{1}}=0.028$$a_{\Delta_{2}}=0$ and $a_{\Delta_{3}}=0.056$. By selecting $\tau_{\text{miati}}^{s}=0.167\ ms$, $\mu=0.47a_{s}\underline{a}_{U_{s}}$ and $\mu_{1}=0.38a_{s}\underline{a}_{U_{s}}$, we obtain $\epsilon^{*}=0.018$, thus the maximum $\tau_{\text{mati}}^{f}$ is 25 ms. Moreover, we have $\gamma_{v_{1}}=8.3$, $\gamma_{v_{2}}=11.5$ and $\gamma_{\dot{u}_{s}}=13.9$. In this example, we aim for the maximum MATI estimates, as a result, the gain ultimate bound is large, as we discussed in Remark 2. For more details, please refers to the proof of Theorem 1.
5.2 Simulation
Figure 2 shows a simulation of the example in Section 5.1 with $\epsilon=0.018$, $\tau_{\text{mati}}^{s}=0.13$ s, $\tau_{\text{mati}}^{f}=7\text{ ms}\leq\epsilon^{*}T^{*}$ and $v_{1}=v_{2}=0$, for both $u_{s,1}=t$ and $u_{s,2}=0.5t+100$. We observe that the ultimate bound is proportional to $|\dot{u}_{s}|$ and independent of $|u_{s}|$.
Figure 2: Simulation with different control inputs
@@ Appendix A Proof of Theorem 1
Moreover, we can always find $\gamma_{v_{\ell}}\geq 0$ such that $\gamma_{\hat{v}_{\ell}}\|\hat{v}_{\ell}\|_{(t,j)}\leq\gamma_{v_{\ell}}\|{v_{\ell}}\|_{(t,j)}$ for $\ell\in\{1,2\}$. For instance, $\gamma_{v_{\ell}}=\gamma_{\hat{v}_{\ell}}$ in sampled-data system. Finally, we proof Theorem 1 by using equation (A.25) in Weixuan_Journal.

### v2

@@ Abstract
thanks: This work was supported by the Australian Research Council under the Discovery Project DP200101303, the France Australia collaboration project IRP-ARS CNRS and the ANR COMMITS ANR-23-CE25-0005, and by ANID through the FONDECYT Initiation grant 11250715. E-mail corresponding author: weixuanw@student.unimelb.edu.au. A full version of this paper is available on https://arxiv.org/abs/2511.16469, see wang2025observer.
@@ 1 Introduction
In addition to the challenges introduced by multiple time scales, many practical systems also involve internal states that cannot be directly measured. For example, in battery management, the state-of-charge and state-of-health are not directly measured and must instead be estimated. In such scenarios, observers are indispensable tools for estimating the unmeasurable states. Moreover, electrochemical, thermal and aging dynamics of battery evolve on multiple time scales, which naturally lead to the use of multi-time-scale observers (zou2016multi). Extensive research has been conducted on the multi-time-scale observer design for both continuous-time and sampled-data singularly perturbed systems, e.g., lin2010composite; yang2016observer; ferrante2021observer, while several studies also investigated observers for NCS, e.g., postoyan2011framework; wang2017observer. Despite these advances, multi–time–scale observers for SPNCS that specifically handle measurement noise, which frequently occurs in practice, while accounting for time–scale separation and network phenomena such as aperiodic transmissions and scheduling protocols remain unexplored.
@@ 4 Stability analysis
$|\xi(t,j)|_{\mathcal{E}}\leq k|\xi(0,0)|_{\mathcal{E}}e^{-a(t+j)}+\gamma_{v_{1}}\|v_{1}\|_{(t,j)}\\ +\gamma_{v_{2}}\|v_{2}\|_{(t,j)}+\gamma_{\dot{u}_{s}}\|\dot{u}_{s}\|_{(t,j)},$ (25)
@@
Proof: See wang2025observer.
Remark 2
It is important to note that our result, presented in Theorem 1, shows that the ultimate bound depends on the derivative of the input signal and the magnitude of the measurement noise. This result is a special case of the standard DISS introduced in DISS.
@@ 5 Illustrative example
Consider plant (1) with $A_{11}=\left[\begin{smallmatrix}a_{1}&0\\ 0&-a_{5}\end{smallmatrix}\right]$, $A_{12}=\left[\begin{smallmatrix}a_{2}&0\\ 0&0\end{smallmatrix}\right]$, $A_{21}=\left[\begin{smallmatrix}0&0\\ a_{3}&0\end{smallmatrix}\right]$, $A_{22}=\left[\begin{smallmatrix}-a_{2}&0\\ -a_{2}&-a_{4}\end{smallmatrix}\right]$, $C_{1s}=\left[\begin{smallmatrix}c_{1}&c_{2}\end{smallmatrix}\right]$, $C_{2s}=\left[\begin{smallmatrix}-1&0\end{smallmatrix}\right]$, $C_{2f}=\left[\begin{smallmatrix}0&-1\end{smallmatrix}\right]$, where $a_{1}=10^{-3}$, $a_{2}=0.37$, $a_{3}=1.1$, $a_{4}=4.9$, $a_{5}=1.2$, $c_{1}=-0.03$ and $c_{2}=1.9$. To simplify the computation, we assume observer gains have the form $L_{1s}=\left[\begin{smallmatrix}n_{1}\\ 0\end{smallmatrix}\right]$, $L_{2s}=\left[\begin{smallmatrix}-n_{2}\\ -n_{2}\end{smallmatrix}\right]$, $L_{1f}=\left[\begin{smallmatrix}0\\ n_{3}\end{smallmatrix}\right]$, and $L_{2f}=\left[\begin{smallmatrix}0\\ 0\end{smallmatrix}\right]$.
@@
Since $u_{s}$, $y_{p_{s}}$ and $y_{p_{f}}$ are scalars, the protocols are given by $h_{s}(\kappa_{s},e_{s})=0$ and $h_{f}(\kappa_{f},e_{f})=0$, which are UGES protocols. Define $W_{\star}(\kappa_{\star},e_{\star})\coloneqq|e_{\star}|$, $s\in\{s,f\}$. Then, Assumption 1 holds with $\underline{a}_{W_{\star}}(s)=\overline{a}_{W_{\star}}(s)=s$, $\lambda_{\star}=0$ and $M_{\star}=1$. Moreover, inequality (18) is satisfied with $L_{s}=1.85$ and $A_{H_{s}}=A_{21}^{s}$, and (20) holds with $L_{f}=0$ and $A_{H_{f}}=A_{21}^{f}$.
@@
By solving the LMI (21) and (23) simultaneously while maximizing $\epsilon^{*}T^{*}$, we obtain the parameters in observer gains to be $n_{1}=0.02$, $n_{2}=0.01$ and $n_{3}=0.18$. Meanwhile, we have $P_{s}=\left[\begin{smallmatrix}0.14&-0.55\\ -0.55&2.8\end{smallmatrix}\right]$, $\gamma_{s}=3.4$, $a_{\rho_{s}}=0.14$, $P_{f}=\left[\begin{smallmatrix}3.5&0.15\\ 0.15&2.7\end{smallmatrix}\right]$ and $\gamma_{f}=1.6$, $a_{\rho_{f}}=2.5$. We select $\tau_{\text{mati}}^{s}=0.15\ s<T(L_{s},\gamma_{s},\lambda_{s})$. Then, with $\lambda_{s}^{*}=0.33$, $\eta_{s1}=0.1$, $\eta_{s2}=0.1$, $\eta_{s3}=0.5$ and $\eta_{s}=\eta_{s1}+\eta_{s2}+\eta_{s3}=0.7$, Lemma 1 is satisfied (see Section A for details). Furthermore, we pick $\lambda_{f}^{*}=0.456$, then $T^{*}=T(L_{f},\gamma_{f},\lambda_{f}^{*})=0.45\ \nicefrac{{s}}{{\epsilon}}$.
@@
For the given $P_{f}$ and $W_{s}$, the minimum possible values for $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$ and $\lambda_{5}$ in (31) are given by $0.0025$, $0.11$, $0.01$, $0.23$ and $0.01$, respectively. Similarly, for the specified $W_{s}$, $P_{s}$, $W_{f}$ and $P_{f}$, it can be shown that $b_{1}=14$, $b_{2}=0.21$, $b_{3}=2.8$, $a_{\Delta_{1}}=0.036$, $a_{\Delta_{2}}=0$ and $a_{\Delta_{3}}=0.056$. By selecting $\tau_{\text{miati}}^{s}=0.149\ ms$, $\mu=0.6a_{s}\underline{a}_{U_{s}}$ and $\mu_{1}=0.4a_{s}\underline{a}_{U_{s}}$, we obtain $\epsilon^{*}=0.016$, thus the maximum $\tau_{\text{mati}}^{f}$ is 7 ms. Moreover, we have $\gamma_{v_{1}}=8.4$, $\gamma_{v_{2}}=8.0$ and $\gamma_{\dot{u}_{s}}=10.0$. In this example, we aim for the maximum MATI estimates, as a result, the gain ultimate bound is large, as we discussed in Remark 3.
To verify the bound (25), we conducted simulations for the example presented above using $\epsilon=0.016$, $\tau_{\text{mati}}^{s}=0.13~\text{s}$, and $\tau_{\text{mati}}^{f}=7~\text{ms}\leq\epsilon^{*}T^{*}$, with initial conditions $x_{p}(0)=(0.1,0.1)$, $z_{p}(0)=(-0.2,-0.2)$, and all other states initialized at the origin. The results are shown in Figures 2 and 3, where two scenarios are considered: (i) a nonzero control input with no measurement noise, and (ii) zero control input subject to measurement noise.
The first scenario, shown in Figure 2, corresponds to $v_{1}=v_{2}=0$, where four different control inputs, namely $u_{s}=t$, $u_{s}=0.5t+100$, $u_{s}=100$ and $u_{s}=0$, are applied. We observe that the input with the larger time derivative (i.e., $u_{s}=t$) results in a larger ultimate bound, even though the input with the smaller derivative (i.e., $u_{s}=0.5t+100$) has a greater magnitude. Furthermore, when $\dot{u}_{s}=0$ (as in the cases $u_{s}=100$ and $u_{s}=0$), $u_{s}$ has no influence on the observer performance. This behaviour is consistent with the bound in (25).
Figure 2: Evolution of $|\xi|_{\mathcal{E}}$ under different $u_{s}$
The second scenario, shown in Figure 3, sets $u_{s}=0$ and introduces measurement noise. Two noise signals, $v_{1}=0.04\sin(25t)$ and $v_{1}=0.02\sin(50t)$, are simulated while $v_{2}=0$. The results show that the measurement noise with larger amplitude yields a larger ultimate bound, even though the time derivatives of the two noise signals have equal magnitude. Overall, the system’s response to both control inputs and measurement noise is consistent with the bound in (25), as illustrated in Figures 2 and 3.
Figure 3: Evolution of $|\xi|_{\mathcal{E}}$ under different $v_{1}$
@@ Appendix A Proof of Theorem 1
Moreover, we can always find $\gamma_{v_{\ell}}\geq 0$ such that $\gamma_{\hat{v}_{\ell}}\|\hat{v}_{\ell}\|_{(t,j)}\leq\gamma_{v_{\ell}}\|{v_{\ell}}\|_{(t,j)}$ for $\ell\in\{1,2\}$. For instance, $\gamma_{v_{\ell}}=\gamma_{\hat{v}_{\ell}}$ in sampled-data system. Finally, we prove Theorem 1 by using equation (A.25) in Weixuan_Journal.